OpenStudy (anonymous):

Help pwease In a certain company, the ratio of programmers to marketers is 3:8, and ratio of customer srvice reps (CSRs) to marketers is 2:3. If there are 27 programmers, there are how many CSRs?

OpenStudy (nincompoop):

any idea how you'd tackle this?

OpenStudy (nincompoop):

ratio is similar to division

OpenStudy (anonymous):

nin

OpenStudy (anonymous):

how do i know what goes on top or bottom in proportions?

OpenStudy (anonymous):

wait how come its different though the marketers are mentioned 2wice and they have a different ratio in both of them

OpenStudy (nincompoop):

let me explain how to decompose this

OpenStudy (anonymous):

k

OpenStudy (nincompoop):

\[\frac{3}{8}=\frac{programmers}{marketers}=\frac{27}{x}\] let me know if this is clear with you before I can move on

OpenStudy (anonymous):

ohhhh because of the ratio oh okay i got it!

OpenStudy (anonymous):

so just to make sure if I'm trying to find the missing x it always has to be a denominator?

OpenStudy (nincompoop):

no it can be in the numerator just like the y

OpenStudy (anonymous):

okay go on

OpenStudy (nincompoop):

read the second to the last post I created, I want you to see this and tell me if you have any objection. We need to be in the same page, and that means I want you to understand it

OpenStudy (nincompoop):

http://www.mathsisfun.com/algebra/proportions.html

OpenStudy (anonymous):

yea i et that the marketers and programers are the same but is it the same just because they're both 3? cuz i feel like if they're a fraction they may not be the same .... but okay go on

OpenStudy (nincompoop):

there is no other way to base it from so we make that assumption

OpenStudy (anonymous):

okay go on

OpenStudy (anonymous):

whoa wtf is going on?

OpenStudy (nincompoop):

http://www.mathsisfun.com/algebra/proportions.html this link has a very short tutorial that will assist you with the very basic concept if you're not following

OpenStudy (anonymous):

y did those get retyped?

OpenStudy (anonymous):

okay ill check it out now

OpenStudy (nincompoop):

I reformatted it so you can see the proportionality including the unknowns

OpenStudy (anonymous):

okay so i get the part and whole part but in this case how do we know what is part and whole how do we know if the 27 is the numerator or denominator in this proportion?

OpenStudy (nincompoop):

do you know the relationship of fraction to percent to decimals?

OpenStudy (anonymous):

yea

OpenStudy (nincompoop):

that is what it is

OpenStudy (anonymous):

wingardium leviosa

OpenStudy (nincompoop):

\[\frac{1}{4} = 0.25 = \frac{25}{100} = 25\%\] look at the pattern 3rd grade math ;)

OpenStudy (anonymous):

dude but they're talkin about programmers and CSR thats not percentages

OpenStudy (nincompoop):

\[\frac{3}{8} = 0.375 = 37.5 \% \] 27 is 37.5% of 72

OpenStudy (anonymous):

so the CSR and Marketers are the whole?

OpenStudy (nincompoop):

per cent means per 100

OpenStudy (anonymous):

DUDE!

OpenStudy (nincompoop):

this is why we move decimal places twice to the right from a decimal format two zeroes, two decimal places

OpenStudy (anonymous):

THEN WHY IS IT